OpenStudy (anonymous):

Find the rate of growth after 8 hours. (Round your answer to three decimal places.)

OpenStudy (ash2326):

@zeenat please post the complete question

OpenStudy (anonymous):

0kay

OpenStudy (anonymous):

A common inhabitant of human intestines is the bacterium Escherichia coli. A cell of this bacterium in a nutrient-broth medium divides into two cells every 20 minutes. The initial population of a culture is 65 cells. (a) Find the relative growth rate. (Assume t is measured in hours.) (b) Find an expression for the number of cells after t hours. (c) Find the number of cells after 8 hours. (d) Find the rate of growth after 8 hours. (Round your answer to three decimal places.) (e) When will the population reach 20,000 cells? (Round your answer to two decimal places.)

OpenStudy (ash2326):

Initial population is 65. It divides into two cells after every 20 minutes so the population doubles every 20 minutes. Do you understand this?

OpenStudy (anonymous):

yeah i do,tnxx

OpenStudy (ash2326):

@zeenat Let the population by y after t minutes. Can you figure out what's the relation between y and initial population 65. ? Remember that after 20 minutes population will be 65*2=130 after 40 minutes it'll be 130*2=260.

OpenStudy (anonymous):

oh yeah, i do, i know what to do now. tnx to you!

OpenStudy (ash2326):

@zeenat tell me what's the relation?

OpenStudy (anonymous):

i tried this t = 65(8^t)In8 which is 8 = 65(8^8)In8... its giving me 2267670593.769, and ts wrong.

OpenStudy (anonymous):

please don't ask me anything, i seriously dont know maths :D

OpenStudy (ash2326):

@zeenat I'll help you but tell me how did you get 8?

OpenStudy (ash2326):

It's doubling so there should be 2 involved, shouldn't be?

OpenStudy (ash2326):

@zeenat I'm here, tell me whenever you need help

OpenStudy (anonymous):

as in how 2 involve? i want you to write it how i will understand. thanks,am really bad in calculus

OpenStudy (ash2326):

Okay We have initial population as 65 and population after t minutes \[\large y=65 \times 2^{\frac{t}{20}}\] so we can see after 20 minutes it's \[\large y=65 \times 2^{\frac{20}{20}}=65\times 2=130\] so this is our equation of the population growth. Now you try the parts if you get stuck, do let me know:)

OpenStudy (anonymous):

i dont know what to do, *sad face*

OpenStudy (ash2326):

@zeenat what don't you understand?

OpenStudy (ash2326):

Let's begin part by part. OK?

OpenStudy (ash2326):

@zeenat I'm waiting for your response?

OpenStudy (anonymous):

okay am here

OpenStudy (anonymous):

answer for a) is In8 b) 65(8^t) c) i inserted 8 in to t.. i got 1090519040

OpenStudy (anonymous):

and part e: a(t) = 20 000 20 000 = 65*8^t 4000/13 = 8^t t = log_8(4000/13) t = 2.755114...

OpenStudy (ash2326):

@zeenat brb

OpenStudy (anonymous):

@ash2326 okay

OpenStudy (ash2326):

I'm back!!

OpenStudy (ash2326):

Oh one thing if t is measured in hours then in 60 minutes the population will increase by 8 , so \[y=65\times 8^t\] that's why there is 8. Do you understand this. Sorry I made a mistake

OpenStudy (anonymous):

its okay, yeah i know that.

OpenStudy (ash2326):

Now the rate of increase \[\frac{dy}{dt}=\frac{d}{dt}65 \times 8^t\] So we get \[\frac{dy}{dt}=65 \times 8^t \times \ln 8\] a) this is the solution for part a b) \[y=65 \times 8^{t}\] Can you try the other parts?

OpenStudy (anonymous):

yeah i did all that but, d) is another story

OpenStudy (ash2326):

put t=8 in \[\frac{dy}{dt}=65 \times 8^{t} \times \ln 8\]

OpenStudy (ash2326):

dy/dt at t=8 is your answer

OpenStudy (anonymous):

okay lemme try

OpenStudy (anonymous):

*wheww*

OpenStudy (ash2326):

What's up?

OpenStudy (anonymous):

am so tired, thank God am done

OpenStudy (ash2326):

Good:D Take rest now:D

OpenStudy (anonymous):

lol, Thanks so much!! i don't know how to thank you, hope you understand :D

OpenStudy (ash2326):

Yeah, :D I'm glad to help:D And \[\large Welcome\ To\ Open\ Study:)\]

OpenStudy (anonymous):

THANK YOU

OpenStudy (anonymous):

you will teach me how to do this thing nxt tym...

OpenStudy (ash2326):

Cool:D I'll teach you :)

OpenStudy (anonymous):

okay, see you when i get stuck again,lol